OpenStudy (anonymous):

The Center of Mass of the Sun-Earth System: The mass of the Sun is 329,390 Earth masses and the mean distance from the center of the Sun to the center of the Earth is 1.496 x 10^8 km. Treating the Earth and Sun as particles, with each mass concentrated at the respective geometric center, how far from the center of the Sun is the center of mass of the Earth-Sun system? Compare this distance with the mean radius of the Sun (6.960 x 10^5 km).

OpenStudy (anonymous):

remember the center of mass formula: if there are "n" particles of masses m1, m2, m3 ... at positions x1, x2, x3 ,.... respectively, then the center of mass is at: \[ \bar{x}=\frac{m1\cdot x1+m2\cdot x2+m3\cdot x3+\cdots}{x1+x2+x3+\cdots} \]

OpenStudy (anonymous):

here, you have m1 = mass of sun, m2 = mass of earth 

OpenStudy (anonymous):

ok i used the same equation but my answer is wrong:S